OpenStudy (anonymous):

Simplify: 4^3x^4/4x^2

OpenStudy (anonymous):

For solving these types of problems: Always remember two things: \[x^a \cdot x^b = x^{a+b}\] \[\frac{x^a}{x^b} = x^{a-b}\]

OpenStudy (anonymous):

Thanks..

OpenStudy (anonymous):

Saying Thanks will not solve your Problem but a Try by yours can...